The patient is a 4-year old child who weighs 32 lb. Ordered: Proventil®1tsp PO tid On hand: Proventil® Syrup, 2mg/ 5mL. According to the package insert, for children 2-6 years of age, dosing should be initialed at 0.1 mg/kg o body weight 3 times a day. Is this order safe? YES____________ NO_______ Why? Show results of calculations. If safe determine the amount to administer. Amount to administer if safe______________(See note above) What do you do if it is not safe.__________________

The amount ordered in teaspsoons has to be converted, correct?
Answer #1

Patient’s weight = 32 lb=14.54 kg

According to the package insert, for children 2-6 years of age, dosing should be initialed at 0.1 mg/kg of body weight 3 times a day.

Thus for this patient as per weight dose will be= 14.54 X 0.1=1.45 mg

Ordered Proventil®= 1tsp PO tid =4.92 ml PO tid

On hand Proventil® Syrup= 2mg/ 5mL= 0.4 mg/ml

Thus 1tsp or 4.92 ml will be =?

5 ml contains =2 mg

1 ml will contain=2/5 mg

4.92 ml will contain=(2/5)x 4.92= 1.97 mg

Means ordered dose is 1.97 mg (after converting as instructed)

And as per patient weight patient as per weight safe dose is =1.45 mg

Ordered dose is more than the safe dose hence it is not safe.

Hence it is advised to adjust the dose as per the weight of the patient and under the recommended safe dose as given above.